Lorimer's mother was 24 years old when Lorimer was born. Now Lorimer's mother is 3 times as old as Lorimer. What are their present ages?
Answer by Cromlix(4381)

Hi there,
Lorimer's mother was 24 years old when Lorimer was born.
M = 24 + L .......(1)
Now Lorimer's mother is 3 times as old as Lorimer.
M = 3L ..........(2)
Replace M in Eq(2) with 3L from Eq(1)
M = 24 + L
3L = 24 + L
Collect like terms:
3L - L = 24
2L = 24
L = 12.
Current ages:
Lorimer 12 years old
Mother 36 years old.
Hope this helps :-)